USING SQUARE ROOT PROPERTY SOLVE THE EQUATION (X-3)^2=81
Answer by stanbon(75887)

USING SQUARE ROOT PROPERTY SOLVE THE EQUATION (X-3)^2=81
=========
(x-3)^2 - 81 = 0
-------
x-3 = 9 or x-3 = -9
x = 12 or x = -6
=====================
Cheers,
Stan H.
---------------------